# When Parts are Greater Than Sums: Individual LLM Components Can Outperform Full Models

### v1

@@ 1 Introduction
Rapid changes in large language models (LLMs) have popularized prompting, which guides LLMs to perform tasks with instructions or examples. Notably, in-context learning (ICL; Brown et al., 2020) adapts LLMs to a new task using only a few labeled examples without parameter updates. However, how LLMs react to the in-context examples is sometimes unintuitive Min et al. (2022). Recently, Sclar et al. (2024) find that even for LLMs with instruction tuning Ouyang et al. (2022) and large size, adding a space or newline in prompts can greatly affect accuracy.
@@
We look into the LLM internals to understand what causes the surprising behavior across various ICL settings. Our work stands in contrast to prior studies on ICL, which treat LLMs as black boxes and alter either the input Liu et al. (2022); Bertsch et al. (2024) or output Zhao et al. (2021); Holtzman et al. (2021). We introduce a new view of ICL by decomposing the output of an LLM into the sum of individual contributions of MLPs and attention heads, denoted “components.” Figure 1 reveals three types of curious components: good-performing (blue) that individually perform well or even outperform the full model, bad-performing (red) that perform below chance, and label-biased (green) that predict the same label on the entire test set. We observe these three types on Llama-2-7B, Llama-2-13B Touvron et al. (2023), and Mistral-Instruct-7B Jiang et al. (2023) across 8 tasks.
We study the sensitivity of LLM components to multiple prompts, formed by different demonstrations and templates. We also construct contrast sets of templates—pairs of similar templates that yield large differences in ICL accuracy. Despite large variance in full-model accuracy, we find that component accuracies correlate well across different demonstrations ($r=0.79$ on average) and contrast set templates ($r=0.57$). The top-performing components in contrast set pairs overlap and achieve decent accuracy even when the full model performs near random (Figure 2). In contrast, the component accuracies of two sampled templates are less correlated ($r=0.34$). Further, good-performing components generalize well to out-of-distribution test sets. For instance, the top-1 component for MNLI outperforms the full Llama-2-13B model by $9.1\%$ on MedNLI; Figure 1 also shows that components are transferrable from SST2 to Yelp. We conclude that components are relatively consistent in their behavior across prompts and datasets.
Inspired by our findings, we propose component reweighting. Compared to prior work that selects prompts to improve ICL accuracy Rubin et al. (2022), component reweighting softly selects components by learning weights from the same few-shot examples to scale component activations. Training these weights only involves learning a linear layer—less than a minute on one CPU. Overall, component reweighting better utilizes the same labeled examples, improving over $24$-shot ICL by $6.0\%,2.2\%,5.1\%$ on Llama-2-7B, Llama-2-13B, and Mistral-Instruct-7B, respectively, while enjoying similar inference speed as 4-shot ICL.
@@ 2.2 A Walkthrough of the Decomposition
where $z=[x^{(0)},\mathbf{h}_{1}^{(1)},\dotsc,\mathbf{h}_{n}^{(L)},m^{(1)},\dotsc,m^{(L)}]$ in Eq. 6 and we index every term in the summation with $j$. LN()($\cdot$)( ⋅ ) denotes the final layernorm, specifically, RMSNorm Zhang and Sennrich (2019) for LLMs in our paper (see A.3). In Eq. 6, LN$(\sum_{j}z_{j})=\frac{\sum_{j}z_{j}}{\text{RMS}(\sum_{j}z_{j})}\odot\gamma$, where RMS denotes root mean square, $\odot$ denotes element-wise multiplication, and $\gamma\in\mathbb{R}^{d}$ is the affine parameters. By pre-computing $\hat{\gamma}=\frac{\gamma}{\text{RMS}(\sum_{j}z_{j})}$, we have:
@@ 4.1 Transfer across Prompt Variants
Table 2 summarizes the results on Llama-2-7B. A.6 shows similar findings on other models. (1) The accuracies of the internal components are highly consistent across different choices of demonstrations, having strong correlations and an average of $0.54$ IoU. (2) The components have much weaker agreement across randomly sampled templates, having a near 0 IoU on BoolQ and QQP. (3) Nevertheless, there is agreement between minimally contrastive templates (Cst T), with an average correlation of $0.572$ across tasks, despite contrasting full-model accuracy. For example, Figure 2 demonstrates that full-model accuracy changes dramatically ($39\%$ vs $89\%$) in a minimal pair of templates, but internal components have a high correlation of $0.81$ and the pair shares top-performing components. Combining (2) and (3) suggests components behave similarly on similar templates, but this similarity decreases as the templates diverge.
@@ 4.3 Transfer between Two Opposite Tasks
We conduct a case study of component transferability across instructions using Task069 and Task070 of Super-NaturalInstructions Wang et al. (2022). The instruction for Task 069 asks for correct answers, while Task070 asks for incorrect ones (“pick the one that makes less sense;” Figure 6). The examples for the two tasks are not parallel.
@@ 5.2 Baselines
Liu et al. (2022) improve ICL accuracy by selecting demonstrations from a pool of labeled data for each test example. Here, we select from the given $K$ labeled examples. Following Rubin et al. (2022), we use SBERT Reimers and Gurevych (2019) to encode examples into sentence embeddings and select the $K^{\prime}=\{3,4\}$ nearest neighbors under cosine similarity as the demonstrations for each test example.
@@
As LLMs tend to predict a certain class over others, Zhao et al. (2021) reweight the output class probabilities. They use context-free inputs, such as “N/A”, to calibrate the probability distribution. However, Fei et al. (2023); Zhou et al. (2023) find context-free inputs sometimes ineffective, because in-domain context is important for calibration. Thus, we introduce Calib+, which calibrates the original probabilities $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ with a training set of in-distribution labeled examples, $\mathcal{D}_{\text{train}}$. We train the calibration weights $\mathbf{v}\in\mathbb{R}^{|\mathcal{Y}|}$ on $\mathcal{D}_{\text{train}}$ with cross-entropy loss and obtain the calibrated probabilities $\hat{\mathbf{p}}=\text{softmax}(\mathbf{v}\cdot\mathbf{p})$. For direct comparisons, we split the $K$ examples into the same $\mathcal{D}_{\text{demo}}$ and $\mathcal{D}_{\text{train}}$ sets as component reweighting for Calib+, where $|\mathcal{D}_{\text{demo}}|=K^{\prime}$. We include the training details of both methods in A.4.
@@ 7 Related Work
Prior work shows that ICL performance varies greatly across different choices of demonstrations and templates Zhao et al. (2021); Lu et al. (2022); Sclar et al. (2024); Voronov et al. (2024). While several approaches, such as prompt selection Liu et al. (2022); Chang and Jia (2023), template ensemble Voronov et al. (2024), and many-shot ICL Agarwal et al. (2024), address this issue, they treat LLMs like black boxes without studying the internals. Besides, they usually greatly increase inference time or require a large set of labeled data, which deviates from true few-shot learning Perez et al. (2021). In comparison, our paper studies this problem by looking inside the LLMs. Rather than selecting prompts, we select components in a soft, learnable way. Our method only requires $\{12,24\}$ examples and has little computation overhead over $4$-shot ICL at inference.
@@
Components interpretation aims to study the function of different components in a trained model Csordás et al. (2021); Elhage et al. (2021); Shah et al. (2024), where components considered in prior work include neurons Radford et al. (2017); Gurnee et al. (2023), attention head Voita et al. (2019); Olsson et al. (2022), and MLPs Geva et al. (2021). To analyze the components, pruning Karnin (1990), probing Alain and Bengio (2017), and early decoding nostalgebraist (2020); Geva et al. (2022) are widely used techniques. For example, Michel et al. (2019) prune away a large percentage of attention heads in Transformer models and show that only a few heads are critical to performance at test time. Li et al. (2024) train a linear probe on every attention head to discover the “truthful heads” inside LLMs. Early decoding interprets the investigated components in the textual space by projecting them through the output embedding matrix. Our model decomposition method is based on early decoding and we share some similarities with prior work Yu et al. (2023); Merullo et al. (2024); Liu et al. (2023); Wang et al. (2023), especially in discovering individual components that perform well. Our contributions lie in providing a new view of ICL via decomposition, revealing the transferability of components across diverse ICL settings.
@@ 9 Limitations
Our component reweighting method requires a small set of labeled data $\mathcal{D}_{\text{train}}$ to train the component weights $w$. However, we believe it is not unreasonable to have at least $K=12$ labeled examples in total and we compare with baselines using the same $K$ examples. On the other hand, we do not compare with fine-tuning-based baselines, such as LM-BFFGao et al. (2021) and LoRA Hu et al. (2021), because they usually require a larger GPU memory and tend to overfit on few-shot examples. Another limitation is that we only experiment with classification tasks for easy evaluation. We leave it for future work to generalize our method to generation tasks by doing decomposition and reweighting at every token during generation.
@@ A.1 Tasks and Templates
Table 7 summarizes all the 13 datasets we use in the paper, where we construct balanced test sets by randomly sampling 2000 examples in each task. We form the prompts by concatenating demonstrations in a randomly shuffled order. To avoid the recency bias Zhao et al. (2021), we keep shuffling the demonstrations until the last two have different labels. For minimally conservative templates (section 4.1), Table 8 compares the contrast sets we construct on Llama-2-7B. For our case study on Task069 and Task070, we sample 3 templates from Sclar et al. (2024). In section 4.3, we calculate the correlation between component accuracies of every pair of runs from the two tasks under a fixed template. The correlation score ($r=-0.6$) is averaged across runs and templates. Figure 6 compare the prompts of Task069 and Task070, which consist of an instruction followed by $K$ templated demonstrations. Originally, the two tasks have $\sim 4\%$ of parallel examples. To make our task transfer challenging, we discard these overlapped examples.
@@ A.3 LayerNorms
Figure 4 shows the transformer architecture in GPT2-like models. Because the layernorms inside each block are before MHA and MLP, known as Pre-LN, Eq. 1 has already taken Ln2 and Ln2 into account, and Eq. 8 only has the term for the final layernorm, LN()($\cdot$)( ⋅ ).
Both Llama-2 and Mistral families use RMSNorm Zhang and Sennrich (2019), a layer normalization variant without centering and adding bias terms. Formally, let $x\in\mathbb{R}^{d}$ be the input, the root mean square norm LN$(x)$ is:
@@
A.4 Training Details and Hyperparameters
A.5 Tests of Significance
@@ A.6 More Results on Transferability
Task069 Mishra et al. (2022); Wang et al. (2022) Abductive NLI 2
Task070 Mishra et al. (2022); Wang et al. (2022) Abductive NLI 2
@@
Figure 6: Comparing the prompts of Task069 and Task070. We follow the formats of Sclar et al. (2024) and prepend the task instructions before $K$ demonstrations. We ensure that the two tasks do not have parallel examples to make the transfer experiment (section 4.3) challenging.

### v2

@@ 1 Introduction
The rapid progress in large language models (LLMs) has popularized prompting, which guides LLMs to perform tasks with instructions or examples. Notably, in-context learning (ICL; Brown et al., 2020) adapts LLMs to a new task using only a few labeled examples without parameter updates. However, how LLMs react to the in-context examples is sometimes unintuitive Min et al. (2022b). Recently, Sclar et al. (2024) find that even for LLMs with instruction tuning Ouyang et al. (2022) or large size, adding a space or newline in prompts can greatly affect accuracy.
@@
We look into the LLM internals to understand what causes the surprising behavior across various ICL settings. Our work stands in contrast to prior studies on ICL, which often treat LLMs as black boxes and alter either the input Chen et al. (2023); Bertsch et al. (2024) or output Zhao et al. (2021); Holtzman et al. (2021). We introduce a new view of ICL by decomposing the output of an LLM into the sum of individual contributions of MLPs and attention heads, denoted “components.” Figure 1 reveals three types of curious components: good-performing (blue) that individually perform well or even outperform the full model, bad-performing (red) that perform below chance, and label-biased (green) that predict the same label on the entire test set. We observe these three types on Llama-2-7B, Llama-2-13B Touvron et al. (2023), and Mistral-Instruct-7B Jiang et al. (2023) across 8 tasks.
We study the sensitivity of LLM components to multiple prompts, formed by different demonstrations and templates. We also construct contrast sets of templates—pairs of similar templates that yield large differences in ICL accuracy. Despite large variance in full-model accuracy, we find that component accuracies correlate well across different demonstrations ($r=0.80$ on average) and contrast set templates ($r=0.57$). The top-performing components in contrast set pairs overlap and achieve decent accuracy even when the full model performs near random (Figure 2). In contrast, the component accuracies of two sampled templates are less correlated ($r=0.34$). Further, good-performing components generalize well to out-of-distribution test sets. For instance, the top-1 component for MNLI outperforms the full Llama-2-13B model by $9.1\%$ on MedNLI; Figure 1 also shows that components are transferrable from SST2 to Yelp. We conclude that components are relatively consistent in their behavior across prompts and datasets.
Inspired by our findings, we propose component reweighting. Compared to prior work that selects prompts to improve ICL accuracy Liu et al. (2022b), component reweighting softly selects components by learning weights from the same few-shot examples to scale component activations. Training these weights only involves learning a linear layer—less than a minute on one CPU. Overall, component reweighting better utilizes the same labeled examples, improving over $24$-shot ICL by $6.0\%,2.2\%,5.1\%$ on Llama-2-7B, Llama-2-13B, and Mistral-Instruct-7B, respectively, while enjoying similar inference speed as 4-shot ICL.
@@ 2.2 A Walkthrough of the Decomposition
where $z=[x^{(0)},\tilde{h}_{1}^{(1)},\dotsc,\tilde{h}_{n}^{(L)},m^{(1)},\dotsc,m^{(L)}]$ in Eq. 6 and we index every term in the summation with $j$. LN()($\cdot$)( ⋅ ) denotes the final layernorm, specifically, RMSNorm Zhang and Sennrich (2019) for LLMs in our paper (see A.3). In Eq. 6, LN$(\sum_{j}z_{j})=\frac{\sum_{j}z_{j}}{\text{RMS}(\sum_{j}z_{j})}\odot\gamma$, where RMS denotes root mean square, $\odot$ denotes element-wise multiplication, and $\gamma\in\mathbb{R}^{d}$ is the affine parameters. By pre-computing $\hat{\gamma}=\frac{\gamma}{\text{RMS}(\sum_{j}z_{j})}$, we have:
@@ 4.1 Transfer across Prompt Variants
Table 2 summarizes the results on Llama-2-7B. A.6 shows similar findings on other models. (1) The accuracies of the internal components are highly consistent across different choices of demonstrations, having strong correlations and an average of $0.54$ IoU. (2) The components have much weaker agreement across randomly sampled templates, having a near 0 IoU on BoolQ and QQP. (3) Nevertheless, there is agreement between minimally contrastive templates (Cst T), with an average correlation of $0.57$ across tasks, despite contrasting full-model accuracy. For example, Figure 2 demonstrates that full-model accuracy changes dramatically ($39\%$ vs $89\%$) in a minimal pair of templates, but internal components have a high correlation of $0.81$ and the pair shares top-performing components. Combining (2) and (3) suggests components behave similarly on similar templates, but this similarity decreases as the templates diverge.
@@ 4.3 Transfer between Two Opposite Tasks
We conduct a case study of component transferability across instructions using Task069 and Task070 of Super-NaturalInstructions Wang et al. (2022b). The instruction for Task069 asks for correct answers, while Task070 asks for incorrect ones (“pick the one that makes less sense;” Figure 8). The examples for the two tasks are not parallel.
@@ 5.2 Baselines
Liu et al. (2022b) improve ICL accuracy by selecting demonstrations from a pool of labeled data for each test example. Here, we select from the given $K$ labeled examples. Following Rubin et al. (2022), we use SBERT Reimers and Gurevych (2019) to encode examples into sentence embeddings and select the $K^{\prime}=\{3,4\}$ nearest neighbors under cosine similarity as the demonstrations for each test example.
@@
As LLMs tend to predict a certain class over others, Zhao et al. (2021) reweight the output class probabilities. They use context-free inputs, such as “N/A”, to calibrate the probability distribution. However, Fei et al. (2023); Zhou et al. (2023) find context-free inputs sometimes ineffective, because in-domain context is important for calibration. Thus, we introduce Calib+, which calibrates the original probabilities $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ with a training set of in-distribution labeled examples, $\mathcal{D}_{\text{train}}$. We train the calibration weights $\mathbf{v}\in\mathbb{R}^{|\mathcal{Y}|}$ on $\mathcal{D}_{\text{train}}$ with cross-entropy loss and obtain the calibrated probabilities $\hat{\mathbf{p}}=\text{softmax}(\mathbf{v}\cdot\mathbf{p})$. For direct comparisons, we split the $K$ examples into the same $\mathcal{D}_{\text{demo}}$ and $\mathcal{D}_{\text{train}}$ sets as component reweighting for Calib+, where $|\mathcal{D}_{\text{demo}}|=K^{\prime}$. We include the training details of both methods in A.7.
@@ 7 Related Work
Prior work shows that ICL performance varies greatly across different choices of demonstrations and templates Zhao et al. (2021); Lu et al. (2022); Sclar et al. (2024); Voronov et al. (2024). While several approaches, such as prompt selection Liu et al. (2022b); Chang and Jia (2023); Fu et al. (2023), prompt ensemble Min et al. (2022a); Arora et al. (2023); Voronov et al. (2024), and many-shot ICL Agarwal et al. (2024), address this issue, they treat LLMs like black boxes without studying the internals. Besides, they greatly increase inference time or require a large set of labeled data, which deviates from true few-shot learning Perez et al. (2021). In comparison, our paper studies this problem by looking inside the LLMs. Rather than selecting prompts, we select components in a soft, learnable way. Our method only requires $\{12,24\}$ examples and has little computation overhead over $4$-shot ICL at inference.
@@
Components interpretation studies the function of different components in a trained model Csordás et al. (2021); Elhage et al. (2021); Shah et al. (2024), where components could be neurons Radford et al. (2017); Wang et al. (2022a); Gurnee et al. (2023), attention heads Olsson et al. (2022), and MLPs Geva et al. (2021). To analyze the components, probing Alain and Bengio (2017), ablation Wang et al. (2023a, b); Chang et al. (2024), patching Wang et al. (2023a); Goldowsky-Dill et al. (2023), and early decoding nostalgebraist (2020); Geva et al. (2022) are widely used techniques. For example, Li et al. (2024) train a linear probe on every attention head to discover the truthful heads inside LLMs. Michel et al. (2019); Voita et al. (2019) prune away a large percentage of attention heads and show that only a few are critical to the performance. Hendel et al. (2023); Liu et al. (2023); Merullo et al. (2024b); Todd et al. (2024) view ICL as compressing demonstrations into function vectors, where they remove the demonstrations and modify (patch) the LLM activations at certain layers with the function vectors at test time. Early decoding interprets the investigated components in the textual space by projecting them through the output embedding matrix Geva et al. (2023). Our model decomposition is based on early decoding and we share some similarities with prior work Yu et al. (2023); Wang et al. (2023c); Merullo et al. (2024a), especially in discovering individual components that perform well on a task. Our contributions lie in providing a new view of ICL by decomposition, which reveals the transferability of components across diverse ICL settings.
@@ 9 Limitations
Our component reweighting method requires a small set of labeled data $\mathcal{D}_{\text{train}}$ to train the component weights $w$. However, we believe it is not unreasonable to have at least $K=12$ labeled examples in total and we compare with baselines using the same $K$ examples. On the other hand, we do not compare with fine-tuning-based baselines, such as LM-BFFGao et al. (2021), T-few Liu et al. (2022a), and LoRA Hu et al. (2021), because they usually require a larger GPU memory for training and more sophisticated early stopping criteria to prevent overfitting on few-shot examples. Another limitation is that we only experiment with classification tasks for easy evaluation. We leave it for future work to generalize our method to generation tasks by doing decomposition and reweighting at every token during generation.
@@ A.1 Tasks and Templates
Table 8 summarizes the 13 datasets we use in the paper, where we construct balanced test sets by randomly sampling 2000 examples in each task. We form the prompts by concatenating demonstrations in a randomly shuffled order. To avoid the recency bias Zhao et al. (2021), we keep shuffling the demonstrations until the last two have different labels. For minimally conservative templates (section 4.1), Table 9 compares the contrast sets we construct on Llama-2-7B. For our case study on Task069 and Task070, we sample 3 templates from Sclar et al. (2024). In section 4.3, we calculate the correlation between component accuracies of every pair of runs from the two tasks under a fixed template. The correlation score ($r=-0.6$) is averaged across runs and templates. Figure 8 compare the prompts of Task069 and Task070, which consist of an instruction followed by $K$ templated demonstrations. Originally, the two tasks have $\sim 4\%$ of parallel examples. To make our task transfer challenging, we discard these overlapped examples.
@@ A.3 LayerNorms
Figure 4 shows the transformer architecture in GPT2-like models. Because the layernorms inside each block are before MHA and MLP, known as Pre-LN, Eq. 1 has already taken $Ln1$ and $Ln2$ into account, and Eq. 6 only has the term for the final layernorm, LN()($\cdot$)( ⋅ ).
Both Llama-2 and Mistral model families use RMSNorm Zhang and Sennrich (2019), a layer normalization variant without centering and adding bias terms. Formally, let $x\in\mathbb{R}^{d}$ be the input, the root mean square norm LN$(x)$ is:
@@
A.4 Tests of Significance
@@
SST2 ARC-Easy
Trained Full $75.8_{\hskip 1.42271pt18.1}$ $57.5_{\hskip 1.42271pt14.4}$
Oracle-T1 $91.7_{\hskip 1.42271pt\phantom{1}0.9}$ $54.5_{\hskip 1.42271pt10.1}$
CompRW $90.8_{\hskip 1.42271pt\phantom{1}1.8}$ $79.0_{\hskip 1.42271pt\phantom{1}1.6}$
Random Full $49.7_{\hskip 1.42271pt0.7}$ $25.0_{\hskip 1.42271pt0.5}$
Oracle-T1 $55.2_{\hskip 1.42271pt0.5}$ $26.7_{\hskip 1.42271pt0.2}$
CompRW $51.4_{\hskip 1.42271pt1.7}$ $25.0_{\hskip 1.42271pt0.7}$
Table 6: Comparing the ICL accuracy between pretrained (Up) and randomly initialized (Down) Llama-2-7B. The top-1 component (Oracle-T1) and CompRW perform near random on the untrained model.
A.5 Do Good-Performing Components Exist in Randomly Initialized Models?
Ramanujan et al. (2020) find that untrained subnetworks can perform on par with a ResNet-34 trained on ImageNet. Similarly, Zhang and Bowman (2018); Hewitt and Liang (2019) show that representations of randomly initialized language models yield a strong baseline for probing tasks. In this section, we investigate (1) whether good-performing components still exist in a randomly initialized LLM, and (2) how CompRW method performs using component activations extracted from the randomly initialized LLM.
We run 4-shot ICL with 15 prompts and report the average accuracy and standard deviation. For CompRW, we use the same 4 demonstrations and 20 more examples for reweighting. Table 6 shows that the best-performing component in a randomly initialized Llama-2-7B still performs poorly on SST2 and ARC-Easy. While CompRW has substantial improvement over Full on the pretrained model, it has no effect on the randomly initialized model. We conclude that good-performing components do not exist in randomly initialized LLMs and our CompRW method relies on the pretrained component activations to perform well.
A.6 More Results on Transferability
A.7 Training Details and Hyperparameters
A.8 Implementation Details
We use the model checkpoints on Hugging Face, Llama-2-7b-hf [^4] https://huggingface.co/meta-llama/Llama-2-7b-hf, Llama-2-13b-hf [^5] https://huggingface.co/meta-llama/Llama-2-13b-hf, and Mistral-7B-Instruct-v0.1666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1. We acknowledge that Mistral-7B-Instruct-v0.3 may have better ICL performance than Mistral-7B-Instruct-v0.1; however, it is not supported by TransformerLens [^7] https://github.com/TransformerLensOrg/TransformerLens (the package we use for model decomposition) to date.
@@ A.8 Implementation Details
Task069 Mishra et al. (2022); Wang et al. (2022b) Abductive NLI 2
Task070 Mishra et al. (2022); Wang et al. (2022b) Abductive NLI 2
@@
Figure 8: Comparing the prompts of Task069 and Task070. We apply the templates of Sclar et al. (2024) and prepend the task instructions before $K$ demonstrations. We ensure that the two tasks do not have parallel examples to make the transfer experiment (section 4.3) challenging.